OpenStudy (anonymous):

Need help with a worded math question please help will give Medal and a Fan (:

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

what grade is this

OpenStudy (anonymous):

well im taking algebra 1

OpenStudy (anonymous):

Can i get your grade

OpenStudy (anonymous):

9th grade

OpenStudy (ranga):

f(x) = -x^2 + 4 The parabola in the diagram has to be moved so that (2, 3) becomes the new vertex. To make (2,3) the vertex of the parabola, the graph has to be shifted to the right by how many units and shifted down by how many units?

OpenStudy (anonymous):

2 to the right? 1 down?

OpenStudy (ranga):

Exactly! Whenever a graph is moved "a" units to the RIGHT, the new function can be obtained by replacing x by (x - a) in the old function. Whenever a graph is moved "a" units to the LEFT, the new function can be obtained by replacing x by (x + a) in the old function. Whenever a graph is moved UP by "a" units, add "a" to the right hand side of the old function to get the new function. Whenever a graph is moved DOWN by "a" units, subtract "a" from the right hand side of the old function to get the new function.

OpenStudy (anonymous):

yay i did it right! lol i didnt expect that

OpenStudy (ranga):

Well done!

OpenStudy (anonymous):

wait is f(x) = -x^2 + 4 the new function? lol sorry if it is and i didnt realize it

OpenStudy (ranga):

old function is: -x^2 + 4. This is the function that is graphed in the link. We need to move this function 2 units to the right and 1 unit down. If a graph is moved 2 units to the RIGHT, replace x with (x-2) in the function: New f(x) = -(x-2)^2 + 4 Next move the graph 1 unit down. If a graph is moved 1 unit DOWN, subtract 1 from the right: New f(x) = -(x-2)^2 + 4 - 1 Expand (x-2)^2 and simplify the above to get your new function.

OpenStudy (anonymous):

ahhhh okay thanks (:

OpenStudy (ranga):

You are welcome.